OpenStudy (anonymous):

Find a solution to the following system of equations. 2x + 4y = 8 x + 4y = 6

OpenStudy (saifoo.khan):

best way is to eliminate y.

OpenStudy (saifoo.khan):

x = 2, y = 1

OpenStudy (anonymous):

x= 2 y= 1

OpenStudy (anonymous):

x + 4y = 6 => x=6-4y substituting x in the first equation 2(6-4y)+4y=8 12-8y+4y=8 12-8=8y-4y 4=4y => y=1 -->(1) substituting (1) in equation 1 x+4(1) =6 x=6-4 x=2 therefore x=2,y=1